Circular No. 7283

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 U1 LONEOS reports the discovery of a new comet. Additional observa-

tions were reported following posting on the NEO Confirmation Page:

1999 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 18.38023 4^{h}38^{m}39.^{s}47 +8^{o}25′00′′.4 17.3 Ferris
18.40064 4 38 38.50 +8 25 04.1  _{”}
18.42104 4 38 37.37 +8 25 06.7  _{”}
18.93387 4 38 11.62 +8 26 35.8  Pravec
18.94006 4 38 11.30 +8 26 36.4  _{”}
18.96318 4 38 10.13 +8 26 40.8 17.2 Tichý
18.96468 4 38 10.09 +8 26 41.2  _{”}
18.96627 4 38 09.98 +8 26 42.1  _{”}
18.96826 4 38 09.84 +8 26 42.6  _{”}
18.96992 4 38 09.81 +8 26 42.4  _{”}

W. D. Ferris (Lowell Observatory). Measurer B. W. Koehn. 0.59-m LO-

NEOS Schmidt + CCD. P. Pravec and P. Kusnirak (Ondřejov). 0.65-m f/3.6 reflector + CCD. Coma

diameter 20′′; broad tail to the south. M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector + CCD. Coma

diameter 10′′ and faint 1′ tail in p.a. 140^{o}.

SUPERNOVA 1999ei IN ANONYMOUS GALAXY J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by M. Wischnjewsky, on a 20-min unfiltered T-Max 400 film taken by L. Gonzalez on Oct. 8.34 UT with the University of Chile’s Maksutov telescope, of a supernova (m_{pv} ∼ 17.5) located at α = 1^{h}33^{m}52^{s}.30, δ = −28^{o}41′57′′.0 (equinox 2000.0), which is ∼ 8′′ west and 1′′ south of the nucleus of the host galaxy. The supernova was confirmed by P. Candia (Cerro Tololo Interamerican Observatory) from B, V, and I images taken with the Cerro Tololo 0.9-m telescope on Oct. 12.12. Further images obtained on Oct. 13.20 by Hamuy with the same telescope give B = 17.4 and V = 17.0. M. M. Phillips (Carnegie Observatories) obtained a spectrogram (range 380–900 nm; resolution 0.7 nm) of SN 1999ei on Oct. 16.19 using the 2.5-m DuPont telescope at Las Campanas Observatory. The spectrum reveals that the supernova is a type-Ia event, ∼ 1 week past maximum. The redshift (based on Hα and [N II] emission lines in the host galaxy) of the supernova is z = 0.031.

1999 October 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green